### v1

@@ 3.1 The exponential decay of $\mathbf{n}\cdot\mathbf{w}$ and $\nabla\cdot\mathbf{w}$ in (39)
Proof.
[…]
@@
$\begin{array}[]{l}\frac{\mathrm{d}}{\mathrm{d}t}\left(e^{2\nu Ct}\|\nabla\cdot\mathbf{w}\|^{2}\right)=e^{2\nu Ct}\left(2\nu C\|\nabla\cdot\mathbf{w}\|^{2}+\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla\cdot\mathbf{w}\|^{2}\right)\leq 0.\end{array}$
Due to the exponential decay in (46), the convergence of $\mathbf{w}$ to $\mathbf{u}$ should be sufficiently fast for most practical applications. In numerical simulations via GePUP-E, it is suggested to set the initial condition of $\mathbf{w}$ to that of $\mathbf{u}$. The mechanism of exponential decay in Theorem 4 is supposed to suppress divergence residue caused by truncation errors of spatial operators; see Theorem 5.
@@ 3.2 Recovering INSE from GePUP-E
Lemma 6.
With the initial condition (40), the GePUP-E formulation (39) recovers the INSE (1) for all $t\geq t_{0}$.
Proof.
(40) and (46) yield $\nabla\cdot\mathbf{w}(t)=0$. With $\mathbf{w}\cdot\mathbf{n}=0$ in Lemma 5, the BVP (42) reduces to a Laplace equation with homogeneous Neumann conditions, for which $\phi$ being a constant in $\Omega$ is a particular solution. Thus $\nabla\phi=\mathbf{0}$ and $\mathbf{w}=\mathbf{u}$. ∎
Lemma 6 implies the well-posedness of (39e,f). In comparison to (21f), (39f) contains the convection term so that the well-posedness of the Neumann BVP (39e,f) is independent of the boundary condition of $\mathbf{u}$.
Lemma 7.
With the initial condition (40), the Neumann BVP (39e,f) admits a unique solution of $\nabla q$.
Proof.
[…]
@@
where the second step follows from Lemma 6. The divergence theorem gives
@@
Then the rest of the proof follows from Theorem 1. ∎
Theorem 5.
@@
Proof.
The first limit follows from (46). The second limit then follows from arguments similar to those in the proof of Lemma 6. The third limit is a consequence of the second limit. ∎
@@ 3.3 The equivalence of GePUP-E with INSE
Lemma 8.
The GePUP-E formulation (39) is derived from the INSE (1) by splitting the velocity $\mathbf{u}$ as $\mathbf{u}=\mathbf{w}+\nabla\phi$ and demanding that $\mathbf{w}$ satisfy (39b) and (44).
Proof.
@@
(39f) follows from Lemma 5 and the normal components of (39a) on $\partial\Omega$, i.e.,
$\begin{array}[]{l}\mathbf{n}\cdot\nabla q=\mathbf{n}\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}+\nu\Delta\mathbf{w})+\frac{\partial(\mathbf{n}\cdot\mathbf{w})}{\partial t}.\end{array}$
An alternative interpretation of (48) might be illuminating. A given scalar field $\phi$ furnishes a specific generic projection ${\mathcal{P}}_{\phi}\mathbf{u}=\mathbf{u}-\nabla\phi$ that perturbs $\mathbf{u}$ to be non-solenoidal. For any $\phi$, the composite operator ${\mathscr{P}}\circ{\mathcal{P}}_{\phi}$ is the identity on the divergence-free vector space, i.e., $\nabla\cdot\mathbf{u}=0$ implies ${\mathscr{P}}\circ{\mathcal{P}}_{\phi}\mathbf{u}={\mathscr{P}}\mathbf{w}=\mathbf{u}$ for any ${\mathcal{C}}^{1}$ scalar field $\phi$. If the evolution of $\mathbf{w}={\mathcal{P}}_{\phi}\mathbf{u}$ in (48) did not have the exponential decay of $\nabla\cdot\mathbf{w}$, we would have to apply ${\mathscr{P}}$ to $\frac{\partial\mathbf{w}}{\partial t}$ to recover the INSE. However, the exponential decay of $\nabla\cdot\mathbf{w}$ in Theorem 4 leads to the convergence of $\mathbf{w}$ to $\mathbf{u}$ in Lemma 5 and thus there is no need to recover $\mathbf{u}$ from $\mathbf{w}$.
Theorem 6.
GePUP-E in Definition 3 is equivalent to the INSE (1).
Proof.
This follows directly from Lemma 6 and Lemma 8. ∎
@@ 3.4 The monotonic decrease of the total kinetic energy
where the second step follows from the divergence theorem and the divergence-free condition of $\mathbf{v}$ and the last from the no-penetration boundary condition. ∎
As a benefit of Lemma 9 to numerical methods, the accuracy of the computed velocity is largely decoupled from that of the pressure gradient; this orthogonality condition can be enforced to machine precision in the FV formulation.
Definition 4.
@@
Theorem 7.
Suppose the body force $\mathbf{g}$ in (39a) is conservative, i.e., $\mathbf{g}=-\nabla\varphi$ for some scalar field $\varphi$ in $\Omega$. Then the evolution of the kinetic energy in the GePUP-E formulation in Definition 3 is governed by
@@
Proof.
Without loss of generality, we may assume that $\mathbf{g}=\mathbf{0}$ since $\varphi$ can be absorbed into $q$. The inner product of $\mathbf{u}$ and the momentum equation (39a) gives
@@
where the first step follows from the definition of $\mathbf{w}$, the second from the commutativity of $\partial_{t}$ and $\nabla$, the third from (39c,d) and Lemma 9, and the last from (51).
The first RHS term in ()(*)( ∗ ) vanishes because
@@
where the fourth step, in Einstein summation convention, follows from Lemma 2.
The second term in ()(*)( ∗ ) also vanishes because of Lemma 9. The third term is
@@
where the first step follows from the definition of $\mathbf{w}$, the second from the commutativity of $\Delta$ and $\nabla$, the third from (39c,d) and Lemma 9, the penultimate from Green’s formula (35), and the last from the no-slip condition of $\mathbf{u}$, which holds from Lemma 6 and (39b,d). ∎
@@ 3.5 Prominent features of the GePUP-E formulation
GePUP-E is summarized as follows.
  1. [(GEF-1)]
  2. 1.
    The sole evolutionary variable is the non-solenoidal velocity $\mathbf{w}$, with $\mathbf{u}$ determined from $\mathbf{w}$ via (39c,d) and $q$ from $\mathbf{u}$ and $\mathbf{w}$ via (39e,f). This chain of determination $\mathbf{w}\rightarrow\mathbf{u}\rightarrow q$ from Neumann BVPs is instantaneous and has nothing to do with time integration. Therefore, a time integrator in MOL can be employed in a black-box manner.
  3. 2.
    There is no ambiguity on which velocities should be projected and which should not in MOL; this resolves (RSN-1) discussed in Section 1.2.3.
  4. 3.
    Now that the main evolutionary variable $\mathbf{w}$ in (39) is formally non-solenoidal, the Leray-Helmholtz projection ${\mathscr{P}}$ only comes into the system (39) on the RHS. Although still contributing to the local truncation error, the approximation error of a discrete projection to ${\mathscr{P}}$ does not affect numerical stability of MOL; this resolves (RSN-2) in Section 1.2.3.
  5. 4.
    (39) comes with the built-in mechanisms of exponential decays of velocity divergence and total kinetic energy, which are conducive to the design of semi-discrete algorithms that ensure numerical stability and preserve physical structures of incompressible flows, c.f. Theorems 9 and 10.
@@ 4 The GePUP-ES formulation
The SAV approach, as introduced in [24, 25], has been proposed to develop time discretization schemes that are both efficient and stable for gradient flows. This approach was originally designed to create schemes that are linear, decoupled, unconditionally stable, and can achieve up to second-order accuracy. It has also been successfully extended to address the Navier-Stokes equations in [30, 31, 32].
More recently, based on the generalized SAV approach [33], Huang et al. have devised high-order consistent splitting schemes for the Navier-Stokes equations, with periodic boundary conditions in [34] and no-slip boundary conditions in [35].
In this section, we couple the GePUP-E formulation (39) to the SAV approach introduced in [24, 25] to deal with the nonlinear convection term $\mathbf{u}\cdot\nabla\mathbf{u}$ so that unconditionally energy-stable numerical schemes can be constructed.
Definition 5.
The GePUP-ES formulation of the INSE on no-slip domains is
$\frac{\partial\mathbf{w}}{\partial t}$ $=\mathbf{g}-r(t)\mathbf{u}\cdot\nabla\mathbf{u}-\nabla q+\nu\Delta\mathbf{w}$ in $\text{in }\Omega,$ (54a)
$\mathbf{w}\cdot\bm{\tau}$ $=0,\ \ \nabla\cdot\mathbf{w}=0$ on $\text{on }\partial\Omega,$ (54b)
@@
$\mathbf{u}$ $=\mathscr{P}\mathbf{w}$ in $\text{in }\Omega,$ (54d)
$\mathbf{u}\cdot\mathbf{n}$ $=0$ on $\text{on }\partial\Omega,$ (54e)
$\Delta q$ $=\nabla\cdot\left(\mathbf{g}-r(t)\mathbf{u}\cdot\nabla\mathbf{u}\right)$ in $\text{in }\Omega,$ (54f)
$\mathbf{n}\cdot\nabla q$ $=\mathbf{n}\cdot(\mathbf{g}-r(t)\mathbf{u}\cdot\nabla\mathbf{u}+\nu\Delta\mathbf{w})+\lambda\mathbf{n}\cdot\mathbf{w}$ on $\text{on }\partial\Omega,$ (54g)
@@
The introduction of the SAV $r(t)\equiv 1$ immediately implies $\frac{\mathrm{d}r}{\mathrm{d}t}=0$. We can define the evolution of $r(t)$ as the ODE in (54c) because, for no-slip conditions, we always have $\left\langle\mathbf{u}\cdot\nabla\mathbf{u},\mathbf{u}\right\rangle=0$. As a newly added evolutionary variable, the SAV $r(t)$ is a double-edged sword. On the one hand, it leads to a tighter coupling between $\mathbf{u}$ and $\mathbf{w}$, which makes it difficult to orchestrate an implicit or semi-implicit RK method as solving a sequence of linear systems; see the discussions in Section 5.2. On the other hand, it preserves the monotonic decrease of the kinetic energy; see Theorems 8 and 10.
Theorem 8.
@@
Proof.
Take inner product with $\mathbf{u}$ in (54a), multiply (54c) by $r(t)$, add up the resulting two equations, and we cancel the integral of the convection term to obtain $\left\langle\frac{\partial\mathbf{w}}{\partial t},\mathbf{u}\right\rangle+r(t)r^{\prime}(t)=\left\langle\mathbf{g}-\nabla q+\nu\Delta\mathbf{w},\mathbf{u}\right\rangle$. The rest of the proof follows from (52), (53), and Lemma 9. ∎
Theorem 4 also holds for GePUP-ES with exactly the same proof. Similarly, the GePUP-ES formulation (54) retains the advantage of GePUP-E that the temporal integration and spatial discretization are completely decoupled. Hence the fourth-order finite-volume discrete operators in [5] can be reused and we will focus on temporal integration hereafter.
@@ 5 Algorithms
Based on the GePUP-ES formulation, we construct numerical algorithms in this section to preserve the monotonic decrease of the modified kinetic energy, the divergence residue, and the magnitude of the normal velocity on the domain boundary.
@@ 5.1 Semi-discrete GePUP-ES-RK schemes
Definition 6.
A GePUP-ES-RK scheme for solving the INSE with no-slip conditions is a semidiscrete algorithm of the form
$\left\{\begin{array}[]{rlcl}\mathbf{w}^{n+1}&=\mathbf{w}^{n}+k\sum_{i=1}^{s}b_{i}\bm{\rho}^{(i)}&&\text{in }\Omega,\\ r^{n+1}&=r^{n}+k\sum_{i=1}^{s}b_{i}I_{cv}(\widetilde{\mathbf{u}}^{(i)},\mathbf{u}^{(i)}),&&\\ \mathbf{u}^{n+1}&=\mathscr{P}\mathbf{w}^{n+1}&&\text{in }\Omega,\\ \mathbf{n}\cdot\mathbf{u}^{n+1}&=0&&\text{on }\partial\Omega,\end{array}\right.$ (55)
where $s$ is the number of stages of the employed RK method $(A,\mathbf{b},\mathbf{c})$, the integral $I_{cv}(\widetilde{\mathbf{u}}^{(i)},\mathbf{u}^{(i)})$ is the same as that in Definition 5, the auxiliary velocity $\widetilde{\mathbf{u}}^{(i)}$ is a suitable explicit approximation to $\mathbf{u}(t^{n}+c_{i}k)$ and
$\Delta q^{(i)}$ $=\nabla\cdot\left(\mathbf{g}^{(i)}-r^{(i)}\widetilde{\mathbf{u}}^{(i)}\cdot\nabla\widetilde{\mathbf{u}}^{(i)}\right)$ in $\text{in }\Omega,$ (56a)
$\mathbf{n}\cdot\nabla q^{(i)}$ $=\mathbf{n}\cdot\left(\mathbf{g}^{(i)}-r^{(i)}\widetilde{\mathbf{u}}^{(i)}\cdot\nabla\widetilde{\mathbf{u}}^{(i)}+\nu\Delta\mathbf{w}^{(i)}\right)+\lambda\mathbf{n}\cdot\mathbf{w}^{(i)}$ on $\text{on }\partial\Omega,$ (56b)

### v2

@@ 3.1 The exponential decay of $\mathbf{n}\cdot\mathbf{w}$ and $\nabla\cdot\mathbf{w}$ in (32)
Proof
(37) follows from the divergence of (32a), (32e), and the commutativity of $\nabla\cdot$ and $\Delta$. Multiply (37) with $\nabla\cdot\mathbf{w}$, integrate over $\Omega$, and we have
@@
$\begin{array}[]{l}\frac{\mathrm{d}}{\mathrm{d}t}\left(e^{2\nu Ct}\|\nabla\cdot\mathbf{w}\|^{2}\right)=e^{2\nu Ct}\left(2\nu C\|\nabla\cdot\mathbf{w}\|^{2}+\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla\cdot\mathbf{w}\|^{2}\right)\leq 0.\end{array}\quad\qed$
Due to the exponential decay in (39), the convergence of $\mathbf{w}$ to $\mathbf{u}$ should be sufficiently fast for most practical applications. In numerical simulations via GePUP-E, it is suggested to set the initial condition of $\mathbf{w}$ to that of $\mathbf{u}$. Then the mechanism of exponential decay in Theorem 3.4 will suppress divergence residue caused by truncation errors of spatial operators; see Theorem 3.7.
@@ 3.2 Recovering INSE from GePUP-E
Lemma 3.5
With the initial condition (33), the GePUP-E formulation (32) recovers the INSE in (1) for all $t\geq t_{0}$.
Proof
(33) and (39) yield $\nabla\cdot\mathbf{w}(t)=0$. With $\mathbf{w}\cdot\mathbf{n}=0$ in Lemma 3.3, the BVP (35) reduces to a Laplace equation with homogeneous Neumann conditions, for which $\phi$ being a constant in $\Omega$ is a particular solution. Thus $\nabla\phi=\mathbf{0}$ and $\mathbf{w}=\mathbf{u}$. ∎
Lemma 3.5 implies the well-posedness of (32e,f). In comparison to (15f), (32f) contains the convection term so that the well-posedness of the Neumann BVP (32e,f) is independent of the boundary condition of $\mathbf{u}$.
Lemma 3.6
With the initial condition (33), the Neumann BVP (32e,f) admits a unique solution of $\nabla q$.
Proof
The divergence theorem and the commutativity of $\nabla\cdot$ and $\Delta$ imply
@@
where the second step follows from Lemma 3.5. The divergence theorem gives
@@
Then the rest of the proof follows from Theorem 2.1. ∎
Theorem 3.7
@@
Proof
The first limit follows from (39). The second limit then follows from arguments similar to those in the proof of Lemma 3.5. The third limit is a consequence of the second limit. ∎
@@ 3.3 The equivalence of GePUP-E with INSE
Lemma 3.8
The GePUP-E formulation (32) is derived from the INSE in (1) by splitting the velocity $\mathbf{u}$ as $\mathbf{u}=\mathbf{w}+\nabla\phi$ and demanding that $\mathbf{w}$ satisfy (32b) and (37).
Proof
@@
(32f) follows from Lemma 3.3 and the normal components of (32a) on $\partial\Omega$, i.e., $\mathbf{n}\cdot\nabla q=\mathbf{n}\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}+\nu\Delta\mathbf{w})+\frac{\partial(\mathbf{n}\cdot\mathbf{w})}{\partial t}$. ∎
An alternative interpretation of (41) might be illuminating. A given scalar field $\phi$ furnishes a specific generic projection ${\mathcal{P}}_{\phi}\mathbf{u}=\mathbf{u}-\nabla\phi$ that perturbs $\mathbf{u}$ to be non-solenoidal. For any $\phi$, the composite operator ${\mathscr{P}}\circ{\mathcal{P}}_{\phi}$ is the identity on the divergence-free vector space, i.e., $\nabla\cdot\mathbf{u}=0$ implies ${\mathscr{P}}\circ{\mathcal{P}}_{\phi}\mathbf{u}={\mathscr{P}}\mathbf{w}=\mathbf{u}$ for any ${\mathcal{C}}^{1}$ scalar field $\phi$. If the evolution of $\mathbf{w}={\mathcal{P}}_{\phi}\mathbf{u}$ in (41) did not have the exponential decay of $\nabla\cdot\mathbf{w}$, we would have to apply $\mathscr{P}$} to $\frac{\partial\mathbf{w}}{\partial t}$ to recover the INSE. However, thanks to the exponential decay of $\nabla\cdot\mathbf{w}$ in Theorem 3.4, $\mathbf{w}$ converges to $\mathbf{u}$ in Lemma 3.7 and thus there is no need to recover $\mathbf{u}$ from $\mathbf{w}$.
Theorem 3.9
GePUP-E in Definition 3.1 is equivalent to the INSE in (1).
Proof
This follows directly from Lemma 3.5 and Lemma 3.8. ∎
@@ 3.4 The monotonic decrease of the total kinetic energy
where the second step follows from the divergence theorem and $\nabla\cdot\mathbf{v}=0$ in $\Omega$ and the last from $\mathbf{v}\cdot\mathbf{n}=0$ on $\partial\Omega$. ∎
As a benefit of Lemma 3.10, the accuracy of the computed velocity is largely decoupled from that of the pressure gradient; this orthogonality condition can be enforced to machine precision in the FV formulation.
Definition 3.11
@@
Theorem 3.12
Suppose the body force $\mathbf{g}$ in (32a) is conservative, i.e., $\mathbf{g}=-\nabla\varphi$ for some scalar field $\varphi$ in $\Omega$. Then the evolution of the kinetic energy in the GePUP-E formulation in Definition 3.1 is governed by
@@
Proof
Since $\varphi$ can be absorbed into $q$, the assumption $\mathbf{g}=\mathbf{0}$ has no loss of generality. The inner product of $\mathbf{u}$ and the momentum equation (32a) give
@@
where the second step follows from the commutativity of $\partial_{t}$ and $\nabla$, the third from (32c,d) and Lemma 3.10, and the last from (44).
The first RHS term in ()(*) vanishes because
@@
where the fourth step, in Einstein summation convention, follows from Lemma 2.6. The second term in ()(*) is also zero due to Lemma 3.10. The third term is
@@
where the first step follows from the definition of $\mathbf{w}$, the second from the commutativity of $\Delta$ and $\nabla$, the third from (32c,d) and Lemma 3.10, the penultimate from Green’s formula (28), and the last from the no-slip condition of $\mathbf{u}$, which holds from Lemma 3.5 and (32b,d). ∎
@@ 3.5 Prominent features of the GePUP-E formulation
  1. (a)
    The sole evolutionary variable is the non-solenoidal velocity $\mathbf{w}$, with $\mathbf{u}$ determined from $\mathbf{w}$ via (32c,d) and $q$ from $\mathbf{u}$ and $\mathbf{w}$ via (32e,f). This chain of determination $\mathbf{w}\rightarrow\mathbf{u}\rightarrow q$ from Neumann BVPs is instantaneous and has nothing to do with time integration. Therefore, a time integrator in MOL can be employed in a black-box manner.
  2. (b)
    There is no ambiguity on which velocities should be projected and which should not in MOL; this resolves (TMR-1) discussed in Section 1.1.3.
  3. (c)
    Now that the main evolutionary variable $\mathbf{w}$ in (32) is formally non-solenoidal, the Leray-Helmholtz projection $\mathscr{P}$} only comes into the system (32) on the RHS. Although still contributing to the local truncation error, the approximation error of a discrete projection to $\mathscr{P}$} does not affect numerical stability of MOL; this resolves (TMR-2) in Section 1.1.3.
  4. (d)
    (32) comes with the built-in mechanisms of exponential decays of velocity divergence and total kinetic energy, which are conducive to the design of semi-discrete algorithms that ensure numerical stability and preserve physical structures of incompressible flows, c.f. Theorems 5.2 and 5.4.
@@ 4 The GePUP-ES formulation
The SAV approach, as introduced in shen18:SAV ; shen19:SAV_Review , has been proposed to develop time discretization schemes that are both efficient and stable for gradient flows. This approach was originally designed to create schemes that are linear, decoupled, unconditionally stable, and up to second-order accurate. It is also successfully extended to address the Navier-Stokes equations in li22:_new_sav_pressure_correction_inse ; lin19:_numer_appx_inses_auxiliary_energy_variable ; li20:_error_analysis_sav_mac .
More recently, based on the generalized SAV approach huang22:_new_class_imex_bdfk_sav , Huang et al. have devised high-order consistent splitting schemes for the Navier-Stokes equations, with periodic boundary conditions in huang21:_stability_error_analysis_higher_order_imex_ns_periodic_bc and no-slip boundary conditions in wu22:_new_class_higher_order_decoupled_inses_rotating .
In this section, we couple the GePUP-E formulation (32) to the SAV approach introduced in shen18:SAV ; shen19:SAV_Review to deal with the nonlinear convection term $\mathbf{u}\cdot\nabla\mathbf{u}$ so that unconditionally energy-stable numerical schemes can be constructed.
Definition 4.1
The GePUP-ES formulation of INSE on no-slip domains is
$\frac{\partial\mathbf{w}}{\partial t}$ $=\mathbf{g}-r(t)\mathbf{u}\cdot\nabla\mathbf{u}-\nabla q+\nu\Delta\mathbf{w}$ $\text{in }\Omega,$ (47a)
$\mathbf{w}\cdot\bm{\tau}$ $=0,\ \ \nabla\cdot\mathbf{w}=0\ \$$\text{on }\partial\Omega,$ (47b)
@@
$\mathbf{u}$ $=\mathscr{P}\mathbf{w}$ $\text{in }\Omega,$ (47d)
$\mathbf{u}\cdot\mathbf{n}$ $=0$ $\text{on }\partial\Omega,$ (47e)
$\Delta q$ $=\nabla\cdot\left(\mathbf{g}-r(t)\mathbf{u}\cdot\nabla\mathbf{u}\right)$ $\text{in }\Omega,$ (47f)
$\mathbf{n}\cdot\nabla q$ $=\mathbf{n}\cdot(\mathbf{g}-r(t)\mathbf{u}\cdot\nabla\mathbf{u}+\nu\Delta\mathbf{w})+\lambda\mathbf{n}\cdot\mathbf{w}$ $\text{on }\partial\Omega,$ (47g)
@@
The introduction of the SAV $r(t)\equiv 1$ immediately implies $\frac{\mathrm{d}r}{\mathrm{d}t}=0$. We can define the evolution of $r(t)$ as the ODE in (47c) because, for no-slip conditions, we always have $\left\langle\mathbf{u}\cdot\nabla\mathbf{u},\mathbf{u}\right\rangle=0$. As a newly added evolutionary variable, the SAV $r(t)$ is a double-edged sword. On the one hand, it leads to a tighter coupling between $\mathbf{u}$ and $\mathbf{w}$, which makes it difficult to orchestrate an implicit or semi-implicit RK method as solving a sequence of linear systems; see the discussions in Section 5.2. On the other hand, it preserves the monotonic decrease of the kinetic energy; see Theorems 4.2 and 5.4.
Theorem 4.2
@@
Proof
Take inner product with $\mathbf{u}$ in (47a), multiply (47c) by $r(t)$, add up the resulting two equations, and we cancel the integral of the convection term to obtain $\left\langle\frac{\partial\mathbf{w}}{\partial t},\mathbf{u}\right\rangle+r(t)r^{\prime}(t)=\left\langle\mathbf{g}-\nabla q+\nu\Delta\mathbf{w},\mathbf{u}\right\rangle$. The rest of the proof follows from (45), (46), and Lemma 3.10. ∎
Theorem 3.4 also holds for GePUP-ES with exactly the same proof. Similarly, the GePUP-ES formulation (47) retains the advantage of GePUP-E that the temporal integration and spatial discretization are completely decoupled. Hence the fourth-order finite-volume discrete operators in zhang16:_GePUP can be reused and we will focus on temporal integration hereafter.
@@ 5 Algorithms
Based on the GePUP-ES formulation, we construct numerical algorithms to preserve the monotonic decrease of the modified kinetic energy, the divergence residue, and the magnitude of the normal velocity on the domain boundary.
@@ 5.1 Semi-discrete GePUP-ES-RK schemes
Definition 5.1
A GePUP-ES-RK scheme for solving INSE with no-slip conditions is a semidiscrete algorithm of the form
$\left\{\begin{array}[]{rlcl}\mathbf{w}^{n+1}&=\mathbf{u}^{n}+k\sum_{i=1}^{s}b_{i}\bm{\rho}^{(i)}&&\text{in }\Omega,\\ r^{n+1}&=r^{n}+k\sum_{i=1}^{s}b_{i}I_{cv}(\widetilde{\mathbf{u}}^{(i)},\mathbf{u}^{(i)}),&&\\ \mathbf{u}^{n+1}&=\mathscr{P}\mathbf{w}^{n+1}&&\text{in }\Omega,\\ \mathbf{n}\cdot\mathbf{u}^{n+1}&=0&&\text{on }\partial\Omega,\end{array}\right$. (48)
where $s$ is the number of stages of the employed RK method $(A,\mathbf{b},\mathbf{c})$, the integral $I_{cv}(\widetilde{\mathbf{u}}^{(i)},\mathbf{u}^{(i)})$ is the same as that in Definition 4.1, the auxiliary velocity $\widetilde{\mathbf{u}}^{(i)}$ is a suitable explicit approximation to $\mathbf{u}(t^{n}+c_{i}k)$ and
$\Delta q^{(i)}$ $=\nabla\cdot\left(\mathbf{g}^{(i)}-r^{(i)}\widetilde{\mathbf{u}}^{(i)}\cdot\nabla\widetilde{\mathbf{u}}^{(i)}\right)$ $\text{in }\Omega,$ (49a)
$\mathbf{n}\cdot\nabla q^{(i)}$ $=\mathbf{n}\cdot\left(\mathbf{g}^{(i)}-r^{(i)}\widetilde{\mathbf{u}}^{(i)}\cdot\nabla\widetilde{\mathbf{u}}^{(i)}+\nu\Delta\mathbf{w}^{(i)}\right)+\lambda\mathbf{n}\cdot\mathbf{w}^{(i)}$ $\text{on }\partial\Omega,$ (49b)